# SimLTD: Simple Supervised and Semi-Supervised Long-Tailed Object Detection

### v2

@@ 1 Introduction
We propose a simple and scalable framework, aptly named SimLTD, to answer this question. We deconstruct the LTD problem into three stages consisting of (i) pre-training on head classes, (ii) transfer learning on tail classes, and (iii) fine-tuning on a small sampled set composed of both head and tail classes. Unlike previous research on multi-stage training [19, 12], our framework allows for the optional use of unlabeled images to further boost LTD. We learn with supplementary unlabeled data in a semi-supervised manner via pseudo-labeling, and therefore do not explicitly rely on any additional instance-level or image-level supervision.
Our work addresses several challenges associated with the head-to-tail model transfer paradigm [18, 39]. For one, we find that the vanilla transfer of data-rich head representations to data-scarce tail classes does not yield sufficient performance improvement because the distribution of head classes is also skewed. Moreover, the naïve application of unlabeled data can aggravate the model’s inductive bias because the unlabeled samples may follow a similar long-tailed distribution. In such case, the trained model will tend to generate more pseudo labels for the head classes, resulting in a larger degree of pseudo-class imbalance. We overcome these obstacles by incorporating data augmentations specifically designed to mitigate class imbalances in both head and tail training stages. Extensive experiments in §4.3 show that stronger pre-trained models on head classes, with and without unlabeled images, transfer well to the tail classes, leading to a desirable solution for long-tailed detection.
@@ 2 Related Work
Leveraging Extra Data for LTD There is an emerging trend of using the abundant weak-supervision of ImageNet labels to solve the LVIS problem. Although LVIS is a detection dataset of natural scenes and ImageNet is a classification dataset of object images, they both share 997 overlapping classes between their vocabularies, the intersection of which provides a rich source of 1.5M extra images to sample additional LVIS instances. To effectively learn from a mixed LVIS-ImageNet dataset with a domain gap, MosaicOS [45] leverages mosaic augmentation [1], whereas Detic [49] and RichSem [27] rely on the CLIP [30] classifier to map the semantic concepts between LVIS and ImageNet targets.
@@ 3.1 The Devil is in the
Figure 4 quantifies the effectiveness of pre-trained representations on detection. For each model, we chop off the detection head consisting of the bounding box classifier and regressor modules learned on the COCO dataset, re-initialize them with random values, and perform transfer learning on the LVIS tail classes. We update only the box classifier and regressor while keeping the rest of the architecture frozen, essentially treating the pre-trained model as a fixed detector. Besides the pre-trained networks, we also assess the Faster R-CNN detector [32] initialized from scratch, except for the pre-trained backbone, as the lower-bound baseline.
@@ 3.2 Step 1: Representation Learning on
We begin with the supervised setting in which we have training data points , where denotes the -th input image and is the -th ground truth annotation containing the box label and coordinates. Let be a learnable detection function training on the image-target pairs to produce the supervised loss :
@@
| SimLTD MixPL [7] | Objects365-unlabeled | DINO | R50 | 447 | 38.3 | 30.8 | 37.5 | 42.6 |
| SimLTD MixPL [7] | COCO-unlabeled2017 | DINO | R50 | 446 | 38.5 | 31.6 | 37.5 | 42.5 |
@@ 3.3 Step 2: Transfer Learning on
We instantiate the tail models for transfer learning by copying the parameters from the pre-trained head models. Let be the supervised tail model and be the semi-supervised counterpart. We train the tail models the same way per Equations 1 and 2, except we update only the classifier and regressor modules, to adapt them to tail classes, while freezing the rest of the networks. The intuition is that pre-trained representations serve as a bootstrapped initializer to train an accurate tail model, according to our analysis in §3.1.
Recall that, unlike common head objects, the tail classes are intrinsically rare so one cannot expect to find abundant occurrences of them in either labeled or unlabeled source. This raises a hurdle for when we wish to train with unlabeled images: there are very few instances of the rare classes in the unlabeled scenes for the teacher model to propose reliable pseudo targets. We sidestep this hurdle by copying and pasting a random subset of rare instances from the labeled training set to the unlabeled images—a new procedure unique to this work. At each training iteration, the teacher model is guaranteed to see an augmented view of sampled rare objects, amid diverse background scenes, which promotes pseudo-label supervision for the student model. Figure 5 illustrates the impact of this technique on pseudo-labeling along with some examples of the augmented unlabeled images, which are subjected to strong photometric and geometric perturbations with cutout [10, 48]. Although the procedure inevitably leads to redundancy and overfitting, our ablation experiments in §4.5 show that it is surprisingly helpful in adapting head representations to the tail models.
@@ 3.4 Step 3: Fine-Tuning on
At this stage, we have two separate models with a shared representation, one optimized on head classes and the other on tail classes. We wish to unify the two models into one for efficient single-pass inference on test samples containing both head and tail categories. Algorithm 1 provides the pseudocode for the merging scheme referenced in Figure 3.
Note that we merge parameters at the detector module and reuse the pre-trained head representations for the rest of the network. We fine-tune the unified detector on composed of instances, or shots, per class sampled from the long-tailed training set. We update only the box classifier and regressor with a reduced learning rate to slowly adapt them to tail classes while preserving the pre-trained accuracy on head classes. We form to include both head and tail classes for exemplar replay to avoid catastrophic forgetting.
@@ 4.5 Ablation Experiments
The design of SimLTD is centered on an intuitive multi-stage training strategy combined with standard data augmentations, without bells and whistles, to overcome the class imbalances in both head and tail datasets. Table 4 shows the contributions of RFS and Copy-Paste in establishing a more robust baseline than was previously possible in the existing literature, by using the simple Faster R-CNN detector with ResNet-50 FPN. We also showcase the viability of multi-stage learning over the naïve single-stage training procedure on the whole long-tailed dataset, which yields markedly worse results.
The Impact of Transfer Learning on As discussed in §3.3, we transfer the pre-trained head representations to optimize on tail classes in Step 2 of our framework. However, this step is optional and may be skipped, because we can initialize the tail classes with random values before fine-tuning. Figure 8 shows that transfer learning on tail classes is a worthwhile task. Across both supervised and semi-supervised settings, transfer learning gives a boost by up to AP and comes with the added bonus of shortening the fine-tuning time by of the required iterations.
@@ 5 Conclusion
We introduced SimLTD, a simple and versatile framework for supervised and semi-supervised long-tailed detection. Standing out from existing work, SimLTD delivers excellent performance and scalability by achieving new state-of-the-art results on the challenging LVIS v1 benchmark—without requiring auxiliary image-level supervision. We hope the practitioner finds utility in our multi-stage training approach and for our work to spur future research aimed at pushing the performance envelope of long-tailed detection.

### v3

@@ 1 Introduction
We propose a simple and scalable framework, aptly named SimLTD, to answer this question. We deconstruct the LTD problem into three stages consisting of (i) pre-training on head classes, (ii) transfer learning on tail classes, and (iii) fine-tuning on a small sample composed of both head and tail classes. Unlike previous research on multi-stage training [19, 12], our framework allows for the optional use of unlabeled images to further boost LTD. We learn with supplementary unlabeled data in a semi-supervised manner via pseudo-labeling, and therefore do not explicitly rely on any additional instance-level or image-level supervision.
Our work addresses several challenges associated with the head-to-tail model transfer paradigm [18, 39]. For one, we find that the vanilla transfer of data-rich head representations to data-scarce tail classes does not yield sufficient performance improvement because the distribution of head classes is also skewed. Moreover, the naïve application of unlabeled data can aggravate the model’s inductive bias because the unlabeled samples may follow a similar long-tailed distribution. In such case, the trained model will tend to generate more pseudo labels for the head classes, resulting in a larger degree of pseudo-class imbalance. We overcome these obstacles by incorporating data augmentations specifically designed to mitigate class imbalances in both head and tail training stages. Extensive experiments in §4.3 reveal that stronger pre-trained models on head classes, with and without unlabeled images, transfer well to the tail classes, leading to a desirable solution for long-tailed detection.
@@ 2 Related Work
Leveraging Extra Data for LTD There is an emerging trend of using the abundant weak supervision of ImageNet labels to solve the LVIS problem. Although LVIS is a detection dataset of natural scenes and ImageNet is a classification dataset of object images, they both share 997 overlapping classes between their vocabularies, the intersection of which provides a rich source of 1.5M extra images to sample additional LVIS instances. To effectively learn from a mixed LVIS-ImageNet dataset with a domain gap, MosaicOS [45] leverages mosaic augmentation [1], whereas Detic [49] and RichSem [27] rely on the CLIP [30] classifier to map the semantic concepts between LVIS and ImageNet targets.
@@ 3.1 The Devil Is in the
Figure 4 quantifies the effectiveness of pre-trained representations on detection. For each model, we chop off the detection head consisting of the bounding box classifier and regressor modules learned on the COCO dataset, re-initialize them with random weights, and perform transfer learning on the LVIS tail classes. We update only the box classifier and regressor while keeping the rest of the architecture frozen, essentially treating the pre-trained model as a fixed detector. Besides the pre-trained networks, we also assess the Faster R-CNN detector [32] initialized from scratch, except for the pre-trained backbone, as the lower-bound baseline.
@@ 3.2 Step 1: Representation Learning on
We begin with the supervised setting in which we have training data points , where denotes the th input image and is the th ground truth annotation containing the box label and coordinates. Let be a learnable detection function training on the image-target pairs to produce the supervised loss :
@@
| SimLTD MixPL [7] | Objects365-unlabeled | DINO | R50 | 447 | 39.1 | 31.5 | 38.5 | 43.1 |
| SimLTD MixPL [7] | COCO-unlabeled2017 | DINO | R50 | 446 | 39.4 | 32.6 | 38.5 | 43.6 |
@@ 3.3 Step 2: Transfer Learning on
We instantiate the tail models for transfer learning by copying the parameters from the pre-trained head models. Let be the supervised tail model and be the semi-supervised counterpart. We train the tail models the same way per Equations 1 and 2, except we update only the classifier and regressor modules, to adapt them to tail classes, while freezing the rest of the networks. The intuition is that pre-trained representations serve as a bootstrapped initializer to learn an accurate tail model, according to our analysis in §3.1.
Recall that, unlike common head objects, the tail classes are intrinsically rare so one cannot expect to find abundant occurrences of them in either labeled or unlabeled source. This raises a hurdle for when we wish to train with unlabeled images: there are very few instances of the rare classes in the unlabeled scenes for the teacher model to propose reliable pseudo targets. We sidestep this hurdle by copying and pasting a random subset of rare instances from the labeled training set to the unlabeled images—a new procedure unique to this work. At each training iteration, the teacher model is guaranteed to see an augmented view of sampled rare objects, amid diverse background scenes, which promotes pseudo-label supervision for the student model. Figure 5 illustrates the impact of this technique on pseudo-labeling along with some examples of the augmented unlabeled images, which are subjected to strong photometric and geometric perturbations with cutout [10, 48]. Although the procedure inevitably leads to redundancy and overfitting, our ablation experiments in §4.5 reveal that it is surprisingly helpful in adapting head representations to the tail models.
@@ 3.4 Step 3: Fine-Tuning on
At this stage, we have two separate models with a shared representation: one optimized on head classes and the other on tail classes. We wish to unify the two models into one for efficient single-pass inference on test samples containing both head and tail classes. Algorithm 1 gives the pseudocode for the merging scheme as referenced in Figure 3.
Note that we merge parameters at the detector module and reuse the pre-trained head representations for the rest of the network. We fine-tune the unified detector on composed of instances, or shots, per class sampled from the training set. We form to include both head and tail categories for exemplar replay. We update only the box classifier and regressor with a reduced learning rate to slowly adapt them to the tail classes, while preserving the pre-trained accuracy to avoid catastrophic forgetting on the head classes.
@@ 4.5 Ablation Experiments
The design of SimLTD is centered on an intuitive multi-stage training strategy combined with standard data augmentations, without bells and whistles, to overcome the class imbalances in both head and tail datasets. Table 4 shows the contributions of RFS and Copy-Paste in establishing a more robust baseline than was previously possible in the existing literature, by using the simple Faster R-CNN detector with ResNet-50 FPN. We also emphasize the advantage of multi-stage learning over the naïve single-stage training procedure on the whole long-tailed dataset, which yields markedly worse results.
The Impact of Transfer Learning on As discussed in §3.3, we transfer the pre-trained head representations to optimize on tail classes in Step 2 of our framework. However, this step is optional and may be skipped, because we can initialize the tail detector with random weights before fine-tuning. Figure 8 shows that transfer learning on tail classes is a worthwhile task. Across both supervised and semi-supervised settings, transfer learning gives a boost by up to AP and comes with the added bonus of shortening the fine-tuning time by of the required iterations.
@@ 5 Conclusion
We introduced SimLTD, a versatile framework for effective supervised and semi-supervised long-tailed detection with unlabeled images. Standing out from existing work, SimLTD delivers excellent performance and scalability to advance the challenging LVIS v1 object detection benchmark—without requiring auxiliary image-level supervision. We hope the practitioner finds utility in our multi-stage training approach and for our work to spur future research aimed at pushing the performance envelope of long-tailed detection.